Question 2 (20 marks)

Sunshine Corporation purchased a new machine on Jan 4, 2018 for $190,000 cash. The machine has a useful life of 10 years or of 25,000 hours. After the useful life the machine will have a residual value of $2,000. The machine was used for 2,500 hours in 2018 and 3,500 hours in 2019.

Required:

  1. Calculate the depreciation expense for 2018 and 2019 under each of the following methods:
  1. Straight-line (3 marks)
  2. Double diminishing-balance (5 marks)
  3. Units-of -production (5 marks)

  1. Calculate the carrying amount at the end of 2019 for the machine when the company is using the Straight-line depreciation method. (3 marks)

  1. Record the journal entry for the cash purchase of the machine on Jan 4, 2018 and for its depreciation expense for the year ended December 31, 2019 under the units-of-production method. (4 marks)
Answer #1

Amounts are in $

(a)

1. Straight line depreciation

Depreciation per year = (190,000-2,000)/10

= 18,800 per year

This remains same for 1st and 2nd year

2. Double diminishing method

Rate = 100%/10 x 2 = 20%

1st year depreciation = 190,000 x 20% = 38,000

2nd year depreciation = (190,000-38,000) x 20% = 30,400

3. Units of production method

1st year depreciation = (190,000-2,000) x (2,500/25,000) = 18,800

2nd year depreciation = (190,000-2,000) x (3,500/25,000) = 26,320

(b)

Carrying amount at the end of 2019 when we use straight line depreciation

= 190,000 - 18,800 - 18,800

= 152,400

(c)

Journal Entry

For cash purchase

On January 4th, 2018

Machinery 190,000

Cash 190,000

(Machinery purchased by paying cash)

Depreciation entry for 2nd year

On December 31st, 2019

Depreciation expense 26,320

Accumulated depreciation 26,320

(Depreciation expense using units of production method accounted)